OpenStudy (anonymous):

7.A dinner plate fits snugly in a square box with perimeter 48 inches. what is the circumference of the plate? "I got 12 phi cm but i don't know if I got it right. 8. Four saucers are part of the same set as the dinner plate in Number 7. Each has a circumference of 15.7 inches. Will they fit, side by side, in the same square box? If so, how many inches will there be between the saucers for padding?

OpenStudy (anonymous):

7.2*\[\pi\]r

OpenStudy (anonymous):

So the answer is 7.2phi?

OpenStudy (anonymous):

48=4*a a=12 R=a/2=6 C=2piR=12pi

OpenStudy (anonymous):

you were correct

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

and what is number 8?

OpenStudy (anonymous):

i don't understand a bit and i want to go sleep and i go to sleep

OpenStudy (anonymous):

 AT and AS are tangents AT=12cm. WHat is the circumference of circle O?